# FractalForensics: Proactive Deepfake Detection and Localization via Fractal Watermarks

### v1

@@ Abstract.
Proactive Deepfake detection via robust watermarks has been raised ever since passive Deepfake detectors encountered challenges in identifying high-quality synthetic images. However, while demonstrating reasonable detection performance, they lack localization functionality and explainability in detection results. Additionally, the unstable robustness of watermarks can significantly affect the detection performance accordingly. In this study, we propose novel fractal watermarks for proactive Deepfake detection and localization, namely FractalForensics. Benefiting from the characteristics of fractals, we devise a parameter-driven watermark generation pipeline that derives fractal-based watermarks and conducts one-way encryption regarding the parameters selected. Subsequently, we propose a semi-fragile watermarking framework for watermark embedding and recovery, trained to be robust against benign image processing operations and fragile when facing Deepfake manipulations in a black-box setting. Meanwhile, we introduce an entry-to-patch strategy that implicitly embeds the watermark matrix entries into image patches at corresponding positions, achieving localization of Deepfake manipulations. Extensive experiments demonstrate satisfactory robustness and fragility of our approach against common image processing operations and Deepfake manipulations, outperforming state-of-the-art semi-fragile watermarking algorithms and passive detectors for Deepfake detection. Furthermore, by highlighting the areas manipulated, our method provides explainability for the proactive Deepfake detection results.
@@ 1. Introduction
The rapid development of Deepfake image manipulation techniques has introduced significant risks to individuals’ cyber privacy and safety, despite offering benefits to society in areas such as entertainment and education (Wang et al., 2024d). While passive Deepfake detection methods (Rössler et al., 2019; Wang et al., 2023; Shiohara and Yamasaki, 2022; Wang and Chow, 2023; Dong et al., 2023; Cao et al., 2022) often face performance bottlenecks when encountering previously unseen generative algorithms, recent studies have begun to explore proactive defenses that intervene prior to image tampering.
@@
Proactive defense using watermarks can be broadly categorized into robust and semi-fragile methods. Although both aim to insert predefined watermarks for image protection and ensure robustness under common image processing operations, the former is designed to maintain robustness even under malicious Deepfake manipulations (Wang et al., 2024a, b), while the latter intentionally allows the watermarks to degrade (Yang et al., 2021; Neekhara et al., 2024). However, existing methods continue to face several key limitations. First, to consistently retain watermark integrity, robust watermarking frameworks usually require training with Deepfake manipulations. Therefore, the robustness tends to significantly degrade when encountering unseen Deepfake manipulation categories [^1] E.g., LampMark (Wang et al., 2024b) has an average of 96.46% watermark recovery rate against face swapping algorithms (trained with) but an average of 73.61% against face reenactment algorithms (unseen). , undermining the reliability of detection results. On the other hand, while allowing fragility when recovering watermarks from Deepfake-manipulated images, the semi-fragile approaches are usually vulnerable to common image processing operations that imitate real-life scenarios. Furthermore, to support forensic investigations in practical Deepfake-related criminal cases, beyond Deepfake detection, the current methods lack explainable localizations of the Deepfake-manipulated area within suspect images. Lastly, since most semi-fragile methods mainly focus on the watermarking frameworks, the watermark confidentiality and ease of use have been barely discussed.
@@
To address the aforementioned issues, we propose FractalForensics, a semi-fragile watermarking approach for proactive Deepfake detection and localization. First, we design a parameter-driven watermark generation pipeline (Figure 1). Specifically, to achieve the localization functionality, we use parameterized fractal curves (Allezaud, 1985) to generate watermarks with spatially meaningful structures and minimal storage overhead. Unlike randomly defined watermark patterns, we devise fractal-based watermarks that are generated following deterministic rules and a small set of parameters, allowing on-demand reconstruction without storing ground-truth layouts. Meanwhile, the recursive and space-filling nature of fractals ensures that watermarks span across most regions of images, which is essential for localizing Deepfake-manipulated areas. On top of the structural backbone, we apply variations to diversify fractal watermark shapes, and to guarantee watermark confidentiality, we further establish a chaotic encryption system (Matthews, 1989) to perform one-way encryption on fractal watermarks, ensuring unpredictability and irreversibility. In general, the entire pipeline is parameter-driven and confidential by design, eliminating the need to store raw watermark data. Then, we construct an end-to-end watermarking framework for embedding and recovery (Figure 2). To preserve spatial correspondence and enable localization, each fractal watermark is represented as a 2-dimensional matrix with 4-bit entries, each encoded as a 4-channel binary tensor. We introduce an entry-to-patch watermark embedding strategy that maps each watermark entry to the image patch at the corresponding position. In particular, for the input image, an image feature extraction module obtains critical image information and a watermark diffusion module expands and refines the input watermark, where outputs of the two modules are passed to the watermark fusion module for watermarked image reconstruction. Thereafter, we construct a decoder for watermark recovery with expected robustness and fragility. In the end, proactive Deepfake detection and localization are accomplished based on the recovered watermarks. The entire framework is trained in a black-box condition without Deepfake manipulations. Extensive experiments further prove the satisfactory Deepfake detection and localization ability of the proposed FractalForensics. The contributions of this work are threefold:
@@
  • Experiments demonstrate promising robustness and fragility of our watermarks regarding common and Deepfake manipulations, outperforming state-of-the-art passive and semi-fragile approaches. Meanwhile, our method generalizes well to unseen datasets and Deepfake manipulations.
@@ 2. Related Work
2.1. Watermarking Approaches for Proactive Deepfake Detection
@@
Although classic passive detectors (Rössler et al., 2019; Wang and Chow, 2023; Shiohara and Yamasaki, 2022; Wang et al., 2023; Cao et al., 2022; Dong et al., 2023) have achieved significant progress since the first occurrence of Deepfake, recent advancements in generative models (Wang et al., 2024c; Xu et al., 2022; Zhao et al., 2023b) have brought strong challenges with high-quality synthetic images. As countermoves, proactive approaches (Huang et al., 2022; Qu et al., 2024; Wang et al., 2025; Wu et al., 2023; Wang et al., 2024a) insert invisible signals into benign images in advance of Deepfake manipulations and actively determine real or fake based on the signals’ functionality afterward.
Robust watermarking frameworks (Zhu et al., 2018; Jia et al., 2021; Ma et al., 2022) are traditionally used for copyright verification and authentication, while recent work assigns semantics to the robust watermarks for Deepfake detection. In 2021, FakeTagger (Wang et al., 2021) raises the general idea of embedding robust and meaningful messages into images for proactive Deepfake detection and provenance. Wang et al. (Wang et al., 2024a) proposed identity perceptual watermarks that are robustly embedded and recovered for the detection of Deepfake face swapping and source tracing based on the consistency in identity information. Later, LampMark (Wang et al., 2024b) exploits the structure-sensitive characteristics and introduces landmark perceptual watermarks, performing proactive Deepfake detection for both face swapping and face reenactment. However, since only trained against the face swapping model, SimSwap (Chen et al., 2020), its watermark robustness significantly drops when evaluated against unseen face reenactment algorithms.
Semi-fragile watermarks are designed to be robust and fragile regarding common image processing operations and Deepfake manipulations, respectively. An early work, FaceGuard (Yang et al., 2021), raises the idea of assigning fragility to watermarks when facing Deepfake manipulations. Later, WaterLo (Beuve et al., 2023) uniformly applies a 1-bit message to the entire benign image and achieves localization for early face swapping algorithms depending on the presence of watermarks in the facial area. Zhao et al. (Zhao et al., 2023a) inserted a watermark into the identity features of the images, and upon modifying facial identities by face swapping manipulations, the image is determined fake based on the absence of the watermark. Similarly, FaceSigns (Neekhara et al., 2024) embeds predefined 128-bit semi-fragile binary watermarks into images via a UNet architecture for Deepfake detection. Recently, EditGuard (Zhang et al., 2024) and OmniGuard (Zhang et al., 2025) treat this task as a united image-bit steganography issue and embed semi-fragile watermarks that accomplish detection and localization regarding AI-generated contents (AIGC) simultaneously. Additionally, some recent studies (Sun et al., 2024, 2022; Wu et al., 2023) propose to utilize robust and semi-fragile watermarks together and achieve proactive detection based on the co-existence of the watermarks, where their semi-fragile watermarks are also worth comparisons and discussions in this study.
@@
Figure 2. Workflow of the proposed FractalForensics. The channel-wise watermark matrix $m^{c}$ is embedded into image $I$ following image feature extraction, watermark diffusion, and watermark fusion modules. To recover the watermark from $I_{\textrm{rec}}$, we construct a decoder. The framework is trained end-to-end. In the testing phase, Deepfake localization is implicitly achieved.
@@ 3.1. Overview
Deepfake manipulates facial areas in full image frames. To protect the faces, we propose FractalForensics, a semi-fragile watermarking framework for proactive Deepfake detection and localization. For a clean facial image $I$ cropped from a full image frame, an encrypted fractal watermark matrix $m$ is generated via a parameter-driven pipeline (Figure 1). This pipeline incorporates user-selected parameters that define the fractal type, structural variation, and chaotic encryption keys. Subsequently, a watermarking framework (Figure 2) is established to invisibly embed $m$ into $I$, consisting of an image feature extraction module, a watermark diffusion module, and a watermark fusion module, to produce the watermarked image $I_{\textrm{rec}}$. A decoder is then constructed to recover the watermark from $I_{\textrm{rec}}$ as $m_{\textrm{rec}}$. By comparing $m_{\textrm{rec}}$ with the original $m$, we achieve proactive Deepfake detection and localization.
@@ 3.2. Watermark Generation Pipeline
Fractals are recursive geometric shapes whose structures are derived based on parameters. Rather than recording all full watermarks in previous studies, we propose a watermark generation pipeline that produces encrypted watermarks depending on the selected parameters. In this study, we adopt the Hilbert curve (Hilbert, 1891) as the selected fractal shape for demonstration. The Hilbert curve of order $n$ defines a continuous path that maps each integer index $i\in[0,2^{2n}-1]$ to a unique coordinate in a $2^{n}\times 2^{n}$ grid. The mapping follows a recursive and self-similar rule that fills the space while preserving spatial locality. Relying on the traversal order defined by the standard Hilbert curve, we construct a 2-dimensional watermark matrix as a representation of the curve.
To ensure diversity in the selected fractal shapes, we apply shape variations toward the standard Hilbert curve by rotations $r\in$0: 0°, 1: 90°, 2: 180°, 3: 270°}, mirroring $m\in$0: None, 1: Top,2: Bottom,3: Left,4: Right,5: Top Left,6: Top Right,7: Bottom Left,8: Bottom Right}, and order modifications $o\in$0: None, 1: Reverse, 2: Zigzag, 3: Cross Flip}, leading to 144 different unique shape variations with respect to the combination of variations. For each curve, there is a start and an end point that record the order and direction [^2] E.g., an $8\times 8$ matrix representing the Hilbert curve has integer entry values from 1 to 64, where 1 is the start and 64 is the end. . To securely protect the curve shape along with the order and direction, we devise a chaotic encryption (Matthews, 1989) system to perform one-way encryption based on the selected parameters. In specific, the logistic map is defined by
@@
where $x_{0}\in[0.1,0.9]^{\mathbb{R}}$ and $a\in[3.7,4.0)^{\mathbb{R}}$ are parameters that determine the chaotic map. To ensure the unpredictability and irreversibility of encrypted watermarks, we let parameters $k\in[100,1000]^{\mathbb{Z}}$ and $d\in[2,20]^{\mathbb{Z}}$ determine the starting value $x_{k}$ of the chaotic sequence and the index of the selected digit of each value. We use the $2^{2n}$ consecutive values at the $d$-th digit starting from $x_{k}$ as the encrypted watermark matrix of the raw entries following the curve shape from the start to the end point. Due to the chaotic characteristic of the encryption system, unless the confidential parameters are known, attempting to recover the original watermark curve by brute force is worse than NP-Hard (van Leeuwen, 1991). For watermarking purposes, since there is more tolerance for binary entries when encoding and decoding, the encrypted watermark entries are transformed into binary formats where each is ultimately a 4-bit entry.
@@ 3.3. Watermarking Framework
We construct a watermarking framework with deep neural networks, aiming to invisibly embed watermark $m$ into image $I$ and perform recovery with robustness and fragility accordingly. To achieve the localization functionality for tampered areas, we introduce an innovative entry-to-patch strategy for watermarking. Specifically, the input image $I$ is implicitly recognized as patches of size $32\times 32$, and the $2^{n}\times 2^{n}$ patches are positionally matched to the watermark matrix $m$ of the same size. To ensure locality preservation such that the watermarks are positionally embedded in the desired locations, we place the 4-bit entries on the channel dimension so that they share the same spatial positions, resulting in the channel-wise binary watermark matrix $m^{c}$ of shape $4\times 2^{n}\times 2^{n}$.
@@
Watermark Diffusion. To enable effective and robust watermark embedding, we expand and refine the binary channel-wise watermark matrix $m^{c}\in\{0,1\}^{4\times 2^{n}\times 2^{n}}$ before fusion for image reconstruction. In specific, the watermark diffusion module first upsamples $m^{c}$ to the same spatial dimension as the input image, where each 4-bit entry is spatially expanded into the corresponding $32\times 32$ patch. Thereafter, the upsampled watermark is fed to a series of a ConvBlock, $N_{\textrm{wtm}}$ SEResBlocks, and a ConvBlock to adaptively refine locally critical watermark features via moderate kernel sizes, ensuring the robustness and invisibility of watermark signals during embedding.
@@
Discriminator. In this work, a discriminator is trained from scratch along with the main FractalForensics framework. Specifically, a ConvBlock and $N_{\textrm{rec}}$ SEResBlocks are applied simply followed by a binary classifier to determine watermarked or raw. The discriminator is used to train the main framework adversarially.
Table 1. Quantitative visual quality evaluation of the watermarked images. Information includes model name, resolution (Res.), watermark size (Num. Bits), PSNR, SSIM, and LPIPS.

### v2

@@ Abstract.
Proactive Deepfake detection via robust watermarks has seen interest ever since passive Deepfake detectors encountered challenges in identifying high-quality synthetic images. However, while demonstrating reasonable detection performance, they lack localization functionality and explainability in detection results. Additionally, the unstable robustness of watermarks can significantly affect the detection performance. In this study, we propose novel fractal watermarks for proactive Deepfake detection and localization, namely FractalForensics. Benefiting from the characteristics of fractals, we devise a parameter-driven watermark generation pipeline that derives fractal-based watermarks and performs one-way encryption of the selected parameters. Subsequently, we propose a semi-fragile watermarking framework for watermark embedding and recovery, trained to be robust against benign image processing operations and fragile when facing Deepfake manipulations in a black-box setting. Moreover, we introduce an entry-to-patch strategy that implicitly embeds the watermark matrix entries into image patches at corresponding positions, achieving localization of Deepfake manipulations. Extensive experiments demonstrate satisfactory robustness and fragility of our approach against common image processing operations and Deepfake manipulations, outperforming state-of-the-art semi-fragile watermarking algorithms and passive detectors for Deepfake detection. Furthermore, by highlighting the areas manipulated, our method provides explainability for the proactive Deepfake detection results [^1] Code will be available at https://github.com/wangty1/fractalforensics..
@@ 1. Introduction
The rapid development of Deepfake image manipulation techniques has introduced significant risks to individuals’ cyber privacy and safety, despite offering benefits to society in areas such as entertainment and education (Wang et al., 2024d). Since passive Deepfake detection methods (Rössler et al., 2019; Luo et al., 2023; Shiohara and Yamasaki, 2022; Kong et al., 2024; Dong et al., 2023; Cao et al., 2022; Luo et al., 2024) often face performance bottlenecks when encountering previously unseen generative algorithms, recent studies have begun to explore proactive defenses that intervene prior to image tampering.
@@
Proactive defense using watermarks can be broadly categorized into robust and semi-fragile methods. Although both aim to insert predefined watermarks for image protection and ensure robustness under common image processing operations, the former is designed to maintain robustness even under malicious Deepfake manipulations (Wang et al., 2024a, b), while the latter intentionally allows the watermarks to degrade (Yang et al., 2021; Neekhara et al., 2024). However, existing methods face several key limitations. First, to consistently retain watermark integrity, robust watermarking frameworks usually require training with Deepfake manipulations. Therefore, the robustness tends to significantly degrade when encountering unseen Deepfake manipulation categories [^2] E.g., LampMark (Wang et al., 2024b) has an average of 96.46% watermark recovery rate against face swapping algorithms (trained with) but an average of 73.61% against face reenactment algorithms (unseen). , undermining the reliability of detection results. On the other hand, while allowing fragility when recovering watermarks from Deepfake-manipulated images, the semi-fragile approaches are usually vulnerable to common image processing operations that imitate real-life scenarios. Furthermore, to support forensic investigations in practical Deepfake-related criminal cases, beyond Deepfake detection, the current methods lack explainable localizations of the Deepfake-manipulated area within suspect images. Lastly, since most semi-fragile methods mainly focus on the watermarking frameworks, the watermark confidentiality and ease of use have been barely discussed.
@@
To address the aforementioned issues, we propose FractalForensics, a semi-fragile watermarking approach for proactive Deepfake detection and localization. First, we design a parameter-driven watermark generation pipeline (Figure 1). Specifically, to achieve the localization functionality, we use parameterized fractal curves (Allezaud, 1985) to generate watermarks with spatially meaningful structures and minimal storage overhead. Unlike randomly defined watermark patterns, we devise fractal-based watermarks that are generated following deterministic rules and a small set of parameters, allowing on-demand reconstruction without storing ground-truth layouts. Meanwhile, the recursive and space-filling nature of fractals ensures that watermarks span across most regions of images, which is essential for localizing Deepfake-manipulated areas. On top of the structural backbone, we apply variations to diversify fractal watermark shapes, and to guarantee watermark confidentiality, we further utilize a chaotic encryption system (Matthews, 1989) to perform one-way encryption on fractal watermarks, ensuring unpredictability and irreversibility. In general, the entire pipeline is parameter-driven and confidential by design, eliminating the need to store raw watermark data.
Then, we construct an end-to-end watermarking framework for embedding and recovery (Figure 2). To preserve spatial correspondence and enable localization, each fractal watermark is represented as a 2-dimensional matrix with 4-bit entries, each encoded as a 4-channel binary tensor. We introduce an entry-to-patch watermark embedding strategy that maps each watermark entry to the image patch at the corresponding position. In particular, for the input image, an image feature extraction module obtains critical image information and a watermark diffusion module expands and refines the input watermark, where outputs of the two modules are passed to the watermark fusion module for watermarked image reconstruction. Thereafter, we construct a decoder for watermark recovery with expected robustness and fragility. In the end, proactive Deepfake detection and localization are accomplished based on the recovered watermarks. The entire framework is trained in a black-box condition without Deepfake manipulations. Extensive experiments further establish the satisfactory Deepfake detection and localization ability of the proposed FractalForensics. The contributions of this work are threefold:
@@
  • Experiments demonstrate promising robustness and fragility of our watermarks under common and Deepfake manipulations, outperforming state-of-the-art passive and semi-fragile approaches. Additionally, our method generalizes well to unseen datasets and Deepfake manipulations.
@@ 2. Related Work
2.1. Watermarks for Deepfake Detection
@@
Although classic passive detectors (Rössler et al., 2019; Wang and Chow, 2023; Shiohara and Yamasaki, 2022; Wang et al., 2023; Kong et al., 2022, 2025) have achieved significant progress since the first occurrence of Deepfake, recent advancements in generative models (Wang et al., 2024c; Xu et al., 2022; Zhao et al., 2023b; Tan et al., 2024) have brought strong challenges with high-quality synthetic images. As countermoves, proactive approaches (Huang et al., 2022; Qu et al., 2024; Wang et al., 2025; Wu et al., 2023; Wang et al., 2024a) insert invisible signals into benign images in advance of Deepfake manipulations and actively determine real or fake based on the signals’ functionality afterwards.
@@
Figure 2. Workflow of the proposed FractalForensics. The channel-wise watermark matrix $w^{c}$ is embedded into image II following image feature extraction, watermark diffusion, and watermark fusion modules. To recover the watermark from $I_{\textrm{rec}}$, we construct a decoder. The framework is trained end-to-end. In the testing phase, Deepfake localization is implicitly achieved.
Robust watermarking frameworks (Zhu et al., 2018; Jia et al., 2021; Ma et al., 2022) are traditionally used for copyright verification and authentication, while recent work assigns semantics to the robust watermarks for Deepfake detection. In 2021, FakeTagger (Wang et al., 2021) raises the general idea of embedding robust and meaningful messages into images for proactive Deepfake detection and provenance. Wang et al. (Wang et al., 2024a) proposed identity perceptual watermarks that are robustly embedded and recovered for the detection of Deepfake face swapping and source tracing based on consistency in identity information. Later, LampMark (Wang et al., 2024b) exploits structure-sensitive characteristics and introduces landmark perceptual watermarks, performing proactive Deepfake detection for both face swapping and face reenactment. However, being trained only against the face swapping model, SimSwap (Chen et al., 2020), its robustness drops significantly when evaluated against unseen face reenactment algorithms.
Semi-fragile watermarks are designed to be robust and fragile regarding common image processing operations and Deepfake manipulations, respectively. An early work, FaceGuard (Yang et al., 2021), raises the idea of assigning fragility to watermarks when facing Deepfake manipulations. Later, Kong et al. (Kong et al., 2022) proposed an innovative dual-branch architecture and achieved state-of-the-art performance in Deepfake localization. WaterLo (Beuve et al., 2023) uniformly applies a 1-bit message to the entire benign image and achieves localization for early face swapping algorithms depending on the presence of watermarks in the facial area. Zhao et al. (Zhao et al., 2023a) inserted a watermark into the identity features of the images, and upon modifying facial identities by face swapping manipulations, the image authenticity is determined based on the absence of the watermark. Similarly, FaceSigns (Neekhara et al., 2024) embeds predefined 128-bit semi-fragile binary watermarks into images via a UNet architecture for Deepfake detection. Recently, EditGuard (Zhang et al., 2024) and OmniGuard (Zhang et al., 2025) treat this task as a joint image-bit steganography issue and embed semi-fragile watermarks that accomplish detection and localization regarding AI-generated contents (AIGC) simultaneously. Additionally, some recent studies (Sun et al., 2024, 2022; Wu et al., 2023) propose to utilize robust and semi-fragile watermarks together and achieve proactive detection based on the co-existence of the watermarks, where their semi-fragile watermarks are also worth comparison and discussion in this study.
@@ 3.1. Overview
Deepfakes manipulate facial areas in full image frames. To protect the faces, we propose FractalForensics, a semi-fragile watermarking framework for proactive Deepfake detection and localization. For a clean facial image $I$ cropped from a full image frame, an encrypted fractal watermark matrix $w$ is generated via a parameter-driven pipeline (Figure 1). This pipeline incorporates user-selected parameters that define the fractal type, structural variation, and chaotic encryption keys. Subsequently, a watermarking framework (Figure 2) is established to invisibly embed $w$ into $I$, consisting of an image feature extraction module, a watermark diffusion module, and a watermark fusion module, to produce the watermarked image $I_{\textrm{rec}}$. A decoder is then constructed to recover the watermark from $I_{\textrm{rec}}$ as $w_{\textrm{rec}}$. By comparing $w_{\textrm{rec}}$ with the original $w$, we achieve proactive Deepfake detection and localization.
@@ 3.2. Watermark Generation Pipeline
Fractals are recursive geometric shapes whose structures are derived based on parameters. Rather than storing all watermarks like in previous studies, we propose a watermark generation pipeline that produces encrypted watermarks depending on the selected parameters. In this study, we use the Hilbert curve (Hilbert, 1891) as the selected fractal shape for demonstration. The Hilbert curve of order $n$ defines a continuous path that maps each integer index $i\in[0,2^{2n}-1]$ to a unique coordinate in a $2^{n}\times 2^{n}$ grid. The mapping follows a recursive and self-similar rule that fills the space while preserving spatial locality. Relying on the traversal order defined by the standard Hilbert curve, we construct a 2-dimensional watermark matrix as a representation of the curve.
To ensure diversity in the selected fractal shapes, we apply shape variations toward the standard Hilbert curve by rotations $r\in$ {0: 0°, 1: 90°, 2: 180°, 3: 270°}, mirroring $m\in$ {0: None, 1: Top, 2: Bottom, 3: Left, 4: Right, 5: Top Left, 6: Top Right, 7: Bottom Left, 8: Bottom Right}, and order modifications $o\in$ {0: None, 1: Reverse, 2: Zigzag, 3: Cross Flip}, leading to 144 different unique shape variations based on the combination of parameters. For each curve, there is a start and an end point that record the order and direction. To securely protect the curve shape along with the order and direction, we utilize a chaotic encryption (Matthews, 1989) system to perform one-way encryption based on the selected parameters. In specific, the logistic map is defined by
@@
where $x_{0}\in[0.1,0.9]^{\mathbb{R}}$ and $a\in[3.7,4.0)^{\mathbb{R}}$ are parameters that determine the chaotic map. To ensure the unpredictability and irreversibility of encrypted watermarks, we let parameters $k\in[100,1000]^{\mathbb{Z}}$ and $d\in[2,20]^{\mathbb{Z}}$ determine the starting value $x_{k}$ of the chaotic sequence and the index of the selected digit of each value. We use the $2^{2n}$ consecutive values at the $d$-th digit starting from $x_{k}$ as the encrypted watermark matrix of the raw entries following the curve shape from the start to the end point. Due to the chaotic characteristic of the encryption system, unless the confidential parameters are known, attempting to recover the original watermark curve by brute force is worse than NP-Hard (van Leeuwen, 1991). For watermarking purposes, since there is more tolerance for binary entries when encoding and decoding, the encrypted watermark entries are transformed into binary formats with 4 bits for each entry.
@@ 3.3. Watermarking Framework
We construct a watermarking framework with deep neural networks, aiming to invisibly embed watermark $w$ into image $I$ and perform recovery with robustness and fragility accordingly. To achieve the localization functionality for tampered areas, we introduce an innovative entry-to-patch strategy for watermarking. Specifically, the input image $I$ is implicitly recognized as patches of size $32\times 32$, and the $2^{n}\times 2^{n}$ patches are positionally matched to the watermark matrix $w$ of the same size. To ensure locality preservation such that the watermarks are positionally embedded in the desired locations, we place the 4-bit entries on the channel dimension so that they share the same spatial positions, resulting in the channel-wise binary watermark matrix $w^{c}$ of shape $4\times 2^{n}\times 2^{n}$.
@@
Watermark Diffusion. To enable effective and robust watermark embedding, we expand and refine the binary channel-wise watermark matrix $w^{c}\in\{0,1\}^{4\times 2^{n}\times 2^{n}}$ before fusion for image reconstruction. In specific, the watermark diffusion module first upsamples $w^{c}$ to the same spatial dimension as the input image, where each 4-bit entry is spatially expanded into the corresponding $32\times 32$ patch. Thereafter, the upsampled watermark is fed to a ConvBlock, $N_{\textrm{wtm}}$ SEResBlocks, and a ConvBlock to adaptively refine locally critical watermark features via moderate kernel sizes, ensuring the robustness and invisibility of watermark signals during embedding.
@@
Discriminator. In this work, a discriminator is trained from scratch along with the main FractalForensics framework. Specifically, a ConvBlock and $N_{\textrm{rec}}$ SEResBlocks are applied simply followed by a binary classifier to determine whether it is watermarked or not. The discriminator is used to train the main framework adversarially.
Table 1. Visual quality evaluation of the watermarked images. Information includes model name, resolution (Res.), watermark size (Num. Bits), PSNR, SSIM, and LPIPS.